# Learning with Local Search MCMC Layers

### v1

@@ 1 Introduction
Models that combine neural networks and combinatorial optimization have recently attracted significant attention [14, 39, 8, 6, 52, 5, 34, 45, 7]. Such models enable the transformation of learned continuous latent representations into structured discrete outputs that satisfy complex constraints. They enrich combinatorial optimization algorithms by providing them with context-dependent features, making decisions more resilient to uncertainty. An important subset of this line of research involves integrating, within a neural network, a linear programming layer of the form:
@@
${\bm{\theta}}\mapsto\operatorname*{\mathsf{argmax}}_{{\bm{y}}\in{\mathcal{Y}}}\leavevmode\nobreak\ \langle{\bm{\theta}},{\bm{y}}\rangle\subseteq\operatorname*{\mathsf{argmax}}_{{\bm{y}}\in\operatorname{\mathsf{conv}}({\mathcal{Y}})}\leavevmode\nobreak\ \langle{\bm{\theta}},{\bm{y}}\rangle,$ (1)
@@
where $\mathcal{Y}$ is a finite set of feasible outputs. In the graphical model and structured prediction literature, this is often called the maximum a posteriori (MAP) problem [53]. The main challenge in using such layers lies in their end-to-end model training. Indeed, as piecewise-constant, discontinuous functions, such layers break the differentiable programming computational graph, and prevent one from backpropagating meaningful gradients from the final output of the model to its parameters.
@@
Many approaches have been proposed to derive relaxations and loss functions for this setting; see Section 2 for a detailed review. As outlined in Table 1, these approaches often differ in the oracle (solver) they assume access to. Some rely on an oracle for solving a continuous relaxation of the above linear program, such as a quadratic or entropy-regularized program. They typically perform a single oracle call per data point. Some other approaches assume access to an oracle for solving the original linear program (i.e., a MAP oracle), but perform multiple oracle calls, for smoothing reasons. Theoretical guarantees for these approaches usually assume an oracle returning exact solutions.
@@
Unfortunately, many problems in operations research are NP-hard in nature, making access to an exact oracle difficult. In contrast, operations research applications often rely on local search heuristics, such as simulated annealing. These heuristics iteratively generate a neighbor of the current solution, and either accept it or reject it based on an acceptance rule. The aim of this work is to provide a theoretically-principled approach for learning with such inexact combinatorial solvers. This is crucial for exploiting widely-used advances in heuristics from the operations research literature as layers in neural networks.
@@
To do so, we propose to leverage the links between neighborhood-based, local search heuristics used to approximately solve combinatorial problems, and Markov chain Monte-Carlo (MCMC) methods used to perform approximate marginal inference in graphical models. These lines of research have evolved separately and the links between them have not been exploited for designing principled machine learning pipelines. We make the following contributions:
  • We propose new differentiable MCMC layers in combinatorial spaces. To do so, we propose to transform problem-specific neighborhood systems that make local search heuristics efficient into proposal distributions to implement MCMC on the combinatorial space of feasible solutions.
@@
  • We show that there exist Fenchel-Young losses [8] (even with a single MCMC iteration) whose stochastic gradients are given by the proposed layer, leading to principled algorithms for both the supervised and unsupervised learning settings, for which we provide a convergence analysis.
@@
  • Using our proposed MCMC layer strongly alleviates the computational bottleneck, even more so with few MCMC iterations. This allows us to handle larger problem instances in the training set, leading to better generalization on large-scale applications [40], which is critical in operations research, as it reduces marginal costs [41].
@@
  • We demonstrate our approach on the EURO Meets NeurIPS 2022 challenge [27], a large-scale, ML-enriched dynamic vehicle routing problem with time windows (DVRPTW), which involves an intractable combinatorial optimization problem. In Appendix A, we also empirically validate the quality of the proposed gradient estimator through abundant experiments.
@@
Differentiable DP (2009, 2018) Entropy Exact marginal DP
@@ 2.1 Problem setup
In this paper, our goal is to learn models with an optimization layer of the form
$\widehat{{\bm{y}}}({\bm{\theta}})\coloneqq\operatorname*{\mathsf{argmax}}\limits_{\bm{y}\in\cal{Y}}\,\langle\bm{\theta},\,\bm{y}\rangle+\varphi(\bm{y}),$ (2)
where ${\bm{\theta}}\in{\mathbb{R}}^{d}$ and $\mathcal{Y}\subset\mathbb{R}^{d}$ is a finite but combinatorially-large set. Eq. 2 is a slight generalization of Eq. 1, where $\varphi:\mathcal{Y}\rightarrow\mathbb{R}$ can be used to incorporate a penalty on the solution. It may no longer be a continuous linear program, as $\varphi$ could potentially be undefined on $\operatorname{\mathsf{conv}}({\mathcal{Y}})$. We focus on settings where this optimization problem is intractable and only heuristic algorithms are available to obtain an approximate solution. Our goal is to integrate NP-hard problems arising in operations research (e.g., routing, scheduling, network design), within a neural network. Unfortunately, many existing approaches simply do not work or lack formal guarantees when used with inexact solvers.
We distinguish between two settings. In the unsupervised setting, our goal will be to learn ${\bm{\theta}}\in{\mathbb{R}}^{d}$ from observations ${\bm{y}}_{1},\dots,{\bm{y}}_{N}\in{\mathbb{R}}^{d}$. In the supervised setting, we will assume that ${\bm{\theta}}=g_{W}({\bm{x}})$ and our goal will be to learn the parameters $W$ from observation pairs $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{N},{\bm{y}}_{N})$.
@@ 2.2 Combinatorial optimization as a layer
Because the function in Eq. 1 is piecewise constant and discontinuous, a frequent strategy consists in introducing regularization in the problem so as to obtain a continuous relaxation. In some cases, we may have access to an oracle for directly solving the regularized problem. For instance, when the unregularized problem can be solved by dynamic programming, its entropic regularization can be computed using a change of semi-ring [31] or by algorithmic smoothing [35]. As another example, interior point solvers can be used to compute a logarithmic barrier regularized solution [33].
We focus on the setting where we only have access to a MAP oracle for solving the original unregularized problem 1, as tackled in several works, or more generally problem 2, as permitted by our work. Frank-Wolfe-like methods can be used to solve the regularized problem using only MAP oracle calls [39, 28]. Another strategy consists in injecting noise perturbations [6] in the oracle. This approach can be shown to be implicitly using regularization. In both cases, a Fenchel-Young loss can be associated, giving a principled way to learn with the optimization layer. However, formal guarantees only hold if the oracle used is exact, and in practice it is typically called multiple times during the forward pass. Our proposal enjoys guarantees even with inexact solvers and a single call.
Regarding differentiation, several strategies are possible. When the approach only needs to differentiate through a (regularized) $\max$, as is the case of Fenchel-Young losses, we can use Danskin’s theorem [13]. When the approach needs to differentiate a (regularized) $\operatorname*{\mathsf{argmax}}$, we can either use autodiff on the unrolled solver iterations or implicit differentiation [3, 1, 9]. Differently, Vlastelica et al. [52] propose to compute gradients via continuous interpolation of the solver.
@@ 2.3 Contrastive divergences
An alternative approach to learning in combinatorial spaces is to use energy-based models (EBMs) [30], which define a distribution over outputs via a parameterized energy function $E_{\bm{\theta}}$:
@@
Therefore, we can perform maximum likelihood estimation (MLE) if we can sample from $\bm{p}_{\bm{\theta}}$, but this is hard both in continuous and combinatorial settings, due to its intractable normalization constant. Contrastive divergences [22, 11, 46] address this by using MCMC to obtain (biased) stochastic gradients. Originally developed for restricted Boltzmann machines with $\mathcal{Y}=\{0,1\}^{d}$ and a Gibbs sampler, they have also been applied in continuous domains via Langevin dynamics [15, 16].
@@
Contrastive divergences rely on MCMC to sample from the current model distribution. Unfortunately, designing an MCMC sampler is usually case by case, and MCMC on discrete domains has received comparatively less attention than continuous domains. Recent efforts adapt continuous techniques, such as Langevin dynamics [55, 47] or gradient-informed proposals [20, 42], to discrete settings. However, these works typically assume simple state space structure, like the hypercube or categorical codebooks, and do not handle complex constraints that are ubiquitous in operations research problems. Sun et al. [48] allow structured state spaces via relaxed constraints in the energy function, yet ignore these structures in the proposal supports. Notably, we emphasize that all these works focus on sampling, not on designing differentiable MCMC layers.
3 Local search based MCMC layers
In this section, we show how to design principled combinatorial layers without relying on exact MAP solvers, by transforming local search heuristics into MCMC algorithms.
@@ 3.1 From local search to MCMC
Local search heuristics [19] iteratively generate a neighbor ${\bm{y}}^{\prime}\in{\mathcal{N}}({\bm{y}^{(k)}})$ of the current solution ${\bm{y}^{(k)}}$, and either accept it or reject it based on an acceptance rule, that depends on the objective function, ${\bm{y}^{(k)}}$ and ${\bm{y}}^{\prime}$. In this context, a neighborhood system $\mathcal{N}$ defines, for each feasible solution ${\bm{y}}\in\mathcal{Y}$, a set of neighbors $\mathcal{N}({\bm{y}})\subseteq\mathcal{Y}$. Neighborhoods are problem-specific, and must respect the structure of the problem, i.e., must maintain solution feasibility. They are typically defined implicitly via a set of allowed moves from ${\bm{y}}$. For instance, Table 2 lists example moves for a vehicle routing problem.
Formally, we denote the neighborhood graph by $G_{\mathcal{N}}\coloneqq(\mathcal{Y},E_{\mathcal{N}})$, where edges are defined by $\mathcal{N}$. We assume the graph is undirected, i.e., $\bm{y}^{\prime}\in\mathcal{N}(\bm{y})$ if and only if $\bm{y}\in\mathcal{N}(\bm{y}^{\prime})$, and without self-loops – i.e., $\bm{y}\notin\mathcal{N}(\bm{y})$. A stochastic neighbor generating function is also provided, in the form of a proposal distribution $q(\bm{y}\,,\cdot\,)$ with support either equal to $\mathcal{N}(\bm{y})$ or $\mathcal{N}(\bm{y})\cup\{\bm{y}\}$.
@@
A well-known example of local search heuristic is simulated annealing (SA) [26]. It is intimately related to Metropolis-Hastings (MH) [21], an instance of a MCMC algorithm. We provide a unified view of both in Algorithm 1.
@@
     Sample a neighbor in $\mathcal{N}(\bm{y}^{(k)})$:$\bm{y}^{\prime}\sim q\left(\bm{y}^{(k)},\,\cdot\,\right)$
@@
     Sample a neighbor in $\mathcal{N}_{s}(\bm{y}^{(k)})$:$\bm{y}^{\prime}\sim q_{s}(\bm{y}^{(k)},\,\cdot)$
@@
The difference lies in the acceptance rule, which incorporates a proposal correction ratio for MH, and in the choice of the sequence of temperatures $(t_{k})_{k\in\mathbb{N}}$. In the case of SA, it is chosen to verify $t_{k}\xrightarrow[]{}0$. In the case of MH, it is such that $t_{k}\equiv t$. In this case, the iterates ${\bm{y}^{(k)}}$ of Algorithm 1 follow a time-homogenous Markov chain on $\mathcal{Y}$, defined by the following transition kernel:
@@
In past work, the link between the two algorithms has primarily been used to show that SA converges to the exact MAP solution in the limit of infinite iterations [36, 17]. Under mild conditions – if the neighborhood graph $G_{\mathcal{N}}$ is connected and the chain is aperiodic – the iterates ${\bm{y}^{(k)}}$ of Algorithm 1 (MH case) converge in distribution to the Gibbs distribution (see Section C.1 for a proof):
@@
Algorithm 1 motivates us to define the combinatorial MCMC layer
@@
where $\bm{\theta}\in\mathbb{R}^{d}$ are logits and $t>0$ is a temperature parameter, defaulting to $t=1$. Naturally, the estimate of ${\widehat{\bm{y}}_{t}}(\bm{\theta})$ returned by Algorithm 1 (MH case) is biased, as the Markov chain cannot perfectly mix in a finite number of iterations, except if it is initialized at ${\pi}_{\bm{\theta},t}$. In Section 4, we will show that this does not hinder the convergence of the proposed learning algorithms. The next proposition, proved in Section C.2, states some useful properties of the proposed layer. {proposition} Let ${\bm{\theta}}\in{\mathbb{R}}^{d}$. We have the following properties:

### v2

@@ 1 Introduction
Models that combine neural networks and combinatorial optimization have recently attracted significant attention (Sadana et al., 2024; Mandi et al., 2024; Donti et al., 2017; Berthet et al., 2020; Bengio et al., 2020; Blondel and Roulet, 2024). They enrich combinatorial optimization algorithms with context-dependent features, making decisions more resilient to uncertainty. An important subset of this line of research integrates, within a neural network, a linear programming layer of the form:
@@
${\bm{\theta}}\mapsto\operatorname*{\mathsf{argmax}}_{{\bm{y}}\in{\mathcal{Y}}}\penalty 10000\ \langle{\bm{\theta}},{\bm{y}}\rangle\subseteq\operatorname*{\mathsf{argmax}}_{{\bm{y}}\in\operatorname{\mathsf{conv}}({\mathcal{Y}})}\penalty 10000\ \langle{\bm{\theta}},{\bm{y}}\rangle,$ (1)
@@
where $\mathcal{Y}$ is a finite set of feasible outputs. Such layers enable the transformation of learned, continuous latent representations into structured, discrete outputs, that satisfy complex constraints. In the graphical models and structured prediction literature, Eq.˜1 is known as the maximum a posteriori (MAP) problem (Wainwright and Jordan, 2008). The main challenge in using such layers lies in end-to-end model training. Indeed, as piecewise-constant functions, they break the differentiable computational graph, and prevent one from backpropagating meaningful gradients for learning.
@@
Many approaches have been proposed to derive relaxations and loss functions for this setting; see Section˜2 for a review. Table˜1 contrasts existing approaches, based on the type of oracle they assume access to. Some rely on an oracle for a regularized version of Eq.˜1, typically performing a single oracle call per data point. Others use a solver for the original problem (i.e., a MAP oracle), requiring multiple calls for smoothing reasons. Theoretical guarantees for these approaches usually assume an oracle returning exact solutions.
@@
Unfortunately, many problems in operations research are NP-hard in nature, making exact oracles impractical. Instead, applications often rely on local search heuristics (e.g., simulated annealing), which iteratively generate and then accept or reject a neighbor of the current solution. We aim to provide a principled approach for learning with such inexact combinatorial solvers. This is crucial for exploiting popular heuristics from the operations research literature as layers in neural networks.
@@
To do so, we propose to leverage links between local search heuristics (used for approximate MAP inference in combinatorial optimization problems) and Markov chain Monte-Carlo (MCMC) methods (used for approximate marginal inference in graphical models). These lines of research have evolved quite separately, and their links remain unexploited for designing principled combinatorial layers.
We make the following contributions:
  • We integrate local search heuristics as differentiable, stochastic layers in neural networks, by converting their neighborhood systems to proposal distributions, turning the local search oracle into a discrete MCMC sampler over the combinatorial set of solutions.
  • We extend our framework to handle local search heuristics that leverage a diversity of neighborhood systems, enabling this class of powerful solvers to be used as a unified MCMC sampler.
@@
  • We show that the proposed layer yields stochastic gradients of a Fenchel-Young loss (Blondel et al., 2020) (even with a single MCMC iteration), leading to principled learning algorithms for conditional and unconditional settings, for which we provide a convergence analysis.
@@
  • The proposed layer reduces the computational bottleneck, especially with few MCMC iterations, enabling larger training instances and better generalization at scale (Parmentier, 2021; 2022).
@@
  • We demonstrate our approach on the EURO Meets NeurIPS 2022 challenge (Kool et al., 2023), a large-scale dynamic vehicle routing problem with time windows, and on binary vector prediction tasks. Abundant additional experiments are included in Appendix˜A of the appendix.
@@
Differentiable DP (2009; 2018) Entropy Exact marginal DP
@@ 2.1 Problem setup
In this paper, our goal is to learn models that incorporate optimization layers of the form:
$\widehat{{\bm{y}}}:{\bm{\theta}}\mapsto\operatorname*{\mathsf{argmax}}\limits_{\bm{y}\in\cal{Y}}\,\langle\bm{\theta},\,\bm{y}\rangle+\varphi(\bm{y}),$ (2)
where $\mathcal{Y}\subset\mathbb{R}^{d}$ is a finite but combinatorially-large set, and $\varphi$ encodes structural costs or preferences on outputs (e.g., routing distances, fixed costs) that do not depend on $\bm{\theta}$} (not to be confused with a regularization term). This formulation therefore extends the standard linear objective in Eq.˜1 by allowing additional problem-specific structure.
We focus on settings where Eq.˜2 is intractable and only heuristic algorithms are available to obtain an approximate solution. Our goal is to integrate NP-hard problems arising in operations research (e.g., routing, scheduling, network design), within a neural network. Unfortunately, many existing approaches lack formal guarantees or simply do not work when used with inexact solvers.
We distinguish between two settings. In the unconditional setting, our goal will be to learn ${\bm{\theta}}\in{\mathbb{R}}^{d}$ from observations ${\bm{y}}_{1},\dots,{\bm{y}}_{N}\in{\mathcal{Y}}$. In the conditional setting, we will assume that ${\bm{\theta}}=g_{W}({\bm{x}})$ and our goal will be to learn the parameters $W$ from observation pairs $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{N},{\bm{y}}_{N})$.
@@ 2.2 Combinatorial optimization as a layer
Since the layer defined in Eq.˜1 is piecewise constant, a frequent strategy consists in introducing regularization in the problem so as to obtain a continuous relaxation. In some cases, we may have access to an oracle for directly solving the regularized problem. For instance, dynamic programming solvers can handle entropic regularization through a change of semi-ring (Li and Eisner, 2009) or algorithmic smoothing (Mensch and Blondel, 2018). As another example, interior point solvers can be used to compute a logarithmic barrier regularized solution (Mandi and Guns, 2020). More recently, McKenzie et al. (2024) handle quadratic regularization by leveraging projection oracles.
We focus on settings where only a MAP oracle is available for the original, unregularized optimization problem. While prior work is often limited to the linear form in Eq.˜1 for the latter, our framework also handles the more general Eq.˜2. Frank-Wolfe-like methods can be used to solve the regularized problem using only MAP oracle calls (Niculae et al., 2018; Krishnan et al., 2015). Another strategy consists in injecting noise perturbations (Berthet et al., 2020) in the oracle, which can be shown to be implicitly using regularization. In both cases, a Fenchel-Young loss can be associated, enabling principled learning. However, formal guarantees require an exact oracle, often called multiple times during the forward pass. Our proposal enjoys guarantees even with inexact solvers and a single call.
Regarding differentiation, several strategies are possible. When the approach only needs to differentiate through a (regularized) $\max$, as is the case of Fenchel-Young losses, we can use Danskin’s theorem (Danskin, 1966). When the approach needs to differentiate a (regularized) $\operatorname*{\mathsf{argmax}}$, we can either use autodiff on the unrolled solver iterations or implicit differentiation (Amos and Kolter, 2017; Agrawal et al., 2019; Blondel et al., 2022). Differently, Vlastelica et al. (2020) propose to compute gradients via continuous interpolation of the solver.
@@ 2.3 Contrastive divergences
An alternative approach to learning in combinatorial spaces is to use energy-based models (EBMs) (Lecun et al., 2006), which define a distribution over outputs via a parameterized energy function $E_{\bm{\theta}}$:
@@
Therefore, we can perform maximum likelihood estimation (MLE) if we can sample from $\bm{p}_{\bm{\theta}}$, but this is hard both in continuous and combinatorial settings, due to its intractable normalization constant. Contrastive divergences (Hinton, 2000; Carreira-Perpiñán and Hinton, 2005; Song and Kingma, 2021) address this by using MCMC to obtain (biased) stochastic gradients. Originally developed for restricted Boltzmann machines with $\mathcal{Y}=\{0,1\}^{d}$ and a Gibbs sampler, they have also been applied in continuous domains via Langevin dynamics (Du and Mordatch, 2020; Du et al., 2021).
@@
Contrastive divergences rely on MCMC to sample the model distribution. Unfortunately, designing MCMC samplers is often case-by-case, and discrete domains have received less attention than continuous ones. Recent efforts adapt continuous techniques, such as Langevin dynamics (Zhang et al., 2022; Sun et al., 2023a) or gradient-informed proposals (Grathwohl et al., 2021; Rhodes and Gutmann, 2022), to discrete settings. However, these works often assume simple state spaces (e.g., the hypercube or categorical codebooks), and do not handle complex constraints ubiquitous in operations research. Sun et al. (2023b) allow structured spaces via relaxed constraints in the energy function, yet ignore these structures in their proposal supports. Notably, we emphasize that all these works focus on sampling, not on designing differentiable MCMC layers.
3 Local search-based MCMC layers
This section introduces our core contribution. We first connect local search heuristics and MCMC methods, then use this link to define a stochastic layer based on a single neighborhood system (Algorithm˜1), and subsequently generalize it to leverage diverse neighborhood systems (Algorithm˜2).
@@ 3.1 From local search to MCMC
Local search heuristics (Gendreau et al., 2010) iteratively generate a neighbor ${\bm{y}}^{\prime}\in{\mathcal{N}}({\bm{y}^{(k)}})$ of the current solution $\bm{y}^{(k)}$}, and either accept it or reject it based on an acceptance rule, that depends on the objective function, $\bm{y}^{(k)}$} and $\bm{y}$$^{\prime}$. In this context, a neighborhood system $\mathcal{N}$ defines, for each feasible solution ${\bm{y}}\in\mathcal{Y}$, a set of neighbors $\mathcal{N}({\bm{y}})\subseteq\mathcal{Y}$.
Neighborhoods are problem-specific, and must respect the structure of the problem, i.e., must maintain solution feasibility. They are typically defined implicitly via a set of allowed moves from $\bm{y}$}. For instance, Table˜2 lists example moves for a vehicle routing problem.
[…]
@@
A well-known example of local search heuristic is simulated annealing (SA) (Kirkpatrick et al., 1983). It is intimately related to Metropolis-Hastings (MH) (Hastings, 1970), an instance of a MCMC algorithm. We provide a unified view of both in Algorithm˜1.
@@
  Sample a neighbor in $\mathcal{N}(\bm{y}^{(k)})$$\bm{y}^{\prime}\sim q\left(\bm{y}^{(k)},\,\cdot\,\right)$
@@
  Sample a neighbor in $\mathcal{N}_{s}(\bm{y}^{(k)})$$\bm{y}^{\prime}\sim q_{s}(\bm{y}^{(k)},\,\cdot)$
@@
The difference lies in the acceptance rule, which incorporates a proposal correction ratio for MH, and in the choice of the sequence of temperatures $(t_{k})_{k\in\mathbb{N}}$. In the case of SA, it is chosen to verify $t_{k}\xrightarrow[]{}0$. In the case of MH, it is such that $t_{k}\equiv t$. In this case, the iterates $\bm{y}^{(k)}$} of Algorithm˜1 follow a time-homogenous Markov chain on $\mathcal{Y}$, defined by the following transition kernel:
@@
In past work, the link between the two algorithms has primarily been used to show that SA converges to the exact MAP solution in the limit of infinite iterations (Mitra et al., 1986; Faigle and Schrader, 1988). Under mild conditions – if the neighborhood graph $G_{\mathcal{N}}$ is connected and the chain is aperiodic, the iterates $({\bm{y}^{(k)}})_{k\in\mathbb{N}}$ of Algorithm˜1 (MH case) converge in distribution to the Gibbs distribution (see Section˜D.1 for a proof):
@@
Algorithm˜1 and this result motivate us to define the combinatorial MCMC layer
@@
where $\bm{\theta}\in\mathbb{R}^{d}$ are logits and $t>0$ is a temperature parameter, defaulting to $t=1$. Computing ${\widehat{\bm{y}}_{t}}({\bm{\theta}})$ is known as the marginal inference problem in the graphical models literature. Naturally, the estimate of ${\widehat{\bm{y}}_{t}}(\bm{\theta})$ returned by Algorithm˜1 (MH case) is biased, as the Markov chain cannot perfectly mix in a finite number of iterations, except if it is initialized at ${\pi}_{\bm{\theta},t}$. In Section˜4, we will show that this does not hinder the convergence of the proposed learning algorithms. The next proposition, proved in Section˜D.2, states some useful properties of the proposed layer. {proposition} Let ${\bm{\theta}}\in{\mathbb{R}}^{d}$. We have the following properties: